# Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise

### v1

@@ Abstract
We consider the discrete-time filtering problem in scenarios where the observation noise is degenerate or low. We focus on the case where the observation equation is a linear function of the state and that additive noise is low or degenerate, however, we place minimal assumptions on the hidden state process. In this scenario we derive new particle filtering (PF) algorithms and, under assumptions, in such a way that as the noise becomes more degenerate a PF which approximates the low noise filtering problem provably inherits the properties of the PF used in the degenerate case. We extend our framework to the case where the hidden states are drawn from a diffusion process. In this scenario we develop new PFs which are robust to both low noise and fine levels of time discretization. We illustrate our algorithms numerically on several examples.
Key Words: Filtering, particle filters, low and degenerate noise, Diffusions, time discretization.
@@
1School of Data Science, The Chinese University of Hong Kong, Shenzen, Shenzhen, CN.
@@ 1 Introduction
The discrete-time filtering problem considers an unobserved Markov chain $\{X_{n}\}_{n\geq 0}$ (the signal or hidden chain) and associated data $\{Y_{n}\}_{n\geq 1}$. The data at any time $k$ are assumed to depend on $\{X_{n}\}_{n\geq 0}$ only by the position $X_{k}$ of the hidden state. The objective of filtering, assuming it is well-defined is to compute expectations with respect to (w.r.t.) the conditional distribution of the signal at each time $n$, given all the data observed up-to that time; we call such a distribution the filter. This problem has numerous applications in statistics, applied mathematics, finance and beyond; we refer the reader to the texts [2, 8, 9] and the references therein for a more complete summary.
@@
In most cases of practical interest, the filter is unavailable analytically and one must resort to numerical approximation of which there are plethora of methods; see [2, 8, 9] for some coverage of the subject. In this article we focus on the approach called particle filters (PF). These are Monte Carlo, simulation-based methods that simulate a collection of $N\in\mathbb{N}$ samples in parallel and undergo what are called sampling and resampling operations. Such methods are very well understood from a mathematical perspective (e.g. [9]) and provide almost surely exact approximations of the filter as $N\rightarrow\infty$. Typically they are effective when the dimension of the hidden state, $d_{x}$ say, is moderate, although there are several attempts to enhance the methodology to some special-case high-dimensional problems (e.g. [14, 15]).
@@
In this article we focus on the case where the observations have low or degenerate noise. For instance one could consider the case:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}$ (1.1)
@@
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}}$, $\Delta\geq 0$, $\{\epsilon_{n}\}_{n\geq 1}$ a sequence of $\mathbb{R}^{d_{y}}$-valued independently and identically distributed (i.i.d.) noises and for technical reasons $1\leq d_{y}<d_{x}$, and $A_{n}$ a sequence of $d_{y}\times d_{x}$ real matrices of full row rank. The low noise case could be interpreted as $\Delta$ small (e.g. $\Delta=10^{-6}$) and the degenerate case with $\Delta=0$. As was noticed by [21], this scenario is important in many applications where it is expected that the signal dynamics is strongly driving the observed data; see [3, 12] for example. From a mathematical perspective, in the case $\Delta=0$ one has to define the filter on a sequence of manifolds $\mathsf{M}_{n}=\{x\in\mathbb{R}^{d_{x}}:y_{n}-A_{n}x=0\}$ and indeed this has been explored, for instance, in [21]. In the context of fitting a single model for example the smoothing distribution (that is of $X_{0},\dots,X_{n}|y_{1},\dots,y_{n}$ with $n$ fixed) sampling on manifolds has been quite well investigated using Markov chain Monte Carlo (MCMC) in several interesting contexts; see for example [5, 7, 10, 11, 19]. The extension to the filtering problem seems only to be considered in [20, 21], although [20] do not consider a manifold that has been defined by the data such as $\mathsf{M}_{n}$. The approach of [21] is to leverage sequential MCMC and as such relies on the afore-mentioned MCMC literature. The methodology there is particularly useful in high-dimensions ($d_{x}$ large) but can be computationally expensive for moderate dimensional problems that the PF itself is well-suited towards and it is this latter direction that we follow in this paper.
The application of PFs for low noise problems seldom works well: it is well-known and illustrated later in this article. In the worst case scenario the results can be very numerically unstable. The main idea of this article is to consider, even in the low noise case, sampling on a sequence of manifolds. For instance, in the context of (1.1) $\widetilde{\mathsf{M}}_{n}=\{(x,\epsilon)\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:y_{n}-A_{n}x-\Delta^{1/2}\epsilon=0\}$. The idea then is if one can define a PF on $\mathsf{M}_{n}$ then the PF on $\widetilde{\mathsf{M}}_{n}$ will ultimately inherit the properties of the PF in the degenerate case ($\Delta=0$). In this paper we explore exactly this issue in the presence of linear observation models such as (1.1) and our contributions can be summarized as follows:
@@
  • We derive new PFs for the low noise and degenerate case for linear observation models. We place minimal assumptions on the hidden process $\{X_{n}\}_{n\geq 0}$.
@@
  • In the scenario of the previous bullet and under assumptions, we prove that the $\mathbb{L}_{r}$-error of the PF in the low noise case will converge as $\Delta\downarrow 0$ to the $\mathbb{L}_{r}$-error of the PF in the degenerate noise case.
@@
  • For additive Gaussian errors in the hidden chain and observations we derive the optimal proposal and prove that the weights of the PF in the low noise case will converge as $\Delta\downarrow 0$ to the weights of the PF in the degenerate noise case.
@@
  • In the case where the hidden dynamics follow a diffusion process, we derive a new PF that is robust not only to low noise but also to fine time discretization error. This appears to be the first PF of its kind.
@@
The significance of this work is not only for filtering. The new PFs in this work could be used for parameter estimation associated to diffusion models, such as was done in [1, 13] using the multilevel Monte Carlo method. We note in terms of the fourth contribution, [6] develop a methodology for the case of hypoelliptic diffusions and degenerate noise (with linear constraints), stating that PFs can be used in their context. However, they do not make the point that this enables the low noise case, which is one of the main messages of this work.
@@ 2.1 State-Space Model
We present a large class of state-space models in the following section. The main idea is to relate what we call the noise parameterization and degenerate noise case in full generality. As we stated in the introduction, however, we are only able to present computational methods for the case of linear observation regimes.
We consider the following state-space model (SSM) for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$ given:
$Y_{n}$ $=H_{n}(X_{n},\epsilon_{n})$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n})$
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$, $(\epsilon_{n},\nu_{n})\in\mathbb{R}^{d_{y}+d_{x}}$ are a sequence of independent noises (independent across $n$ and of each other) with positive and smooth Lebesgue density, $H_{n}:\mathbb{R}^{d_{x}+d_{y}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are a sequence of continuously differentiable functions and whose Jacobian is of full row rank.
@@ 2.2 Natural Parametrization
In the given scenario we can define the filtering densities w.r.t. $d_{x}$-dimensional Lebesgue measure as follows. Denoting the positive conditional Lebesgue densities of $Y_{n}|X_{n}$ and $X_{n}|X_{n-1}$ as $h_{n}(y_{n}|x_{n})$ and $f_{n}(x_{n}|x_{n-1})$ respectively we have that the filter (see e.g. [2] and the references therein) at time 1 is
$\pi_{1}(x_{1})=\frac{h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathbb{R}^{d_{x}}}h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})\lambda_{d_{x}}(dx_{1})}$
where $\lambda_{d_{x}}$ is $d_{x}$-dimensional Lebesgue measure and we supress the dependence of $y_{1}$ in the filter. The filter at any subsequent time point $n\geq 1$ has the following prediction-updating formula:
$\pi_{n}(x_{n})=\frac{h_{n}(y_{n}|x_{n})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})}{\int_{\mathbb{R}^{d_{x}}}h_{n}(y_{n}|x_{1})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})\lambda_{d_{x}}(dx_{n})}.$
In practice one cannot compute the filter and resorts to numerical methods. Consider when $d_{y}=1$ and $\epsilon_{n}\sim\mathcal{N}(0,\sigma_{n}^{2})$ with $\mathcal{N}(0,\sigma_{n}^{2})$ the one-dimensional normal distribution of mean zero and variance $\sigma_{n}^{2}$ very small say $10^{-6}$ we call this a special case of low observational noise. In such cases it can be difficult to design an efficient numerical methods and to help deal with such a scenario we will consider a SSM under a noise-based parameterization which we now detail.

### v2

@@ Abstract
We consider the discrete-time filtering problem in scenarios where the observation noise is low or degenerate. We focus on the case where the observation equation is a linear function of the state and the data involve additive noise. However, we place minimal assumptions on the hidden state process. For such a class of models we derive new particle filters (PFs) with the key property that their performance is robust to the size of the observation noise. As a consequence, the developed PFs are well-defined in the limiting case of degenerate observation noise. Indicatively, we prove (under assumptions) that the PF applied in this low noise setting inherits the properties of the PF used in the degenerate case. We extend our framework to the case where the hidden states are drawn from a diffusion process. In this scenario we develop new PFs which are robust to both low noise and fine levels of time discretization. We illustrate our algorithms numerically on several examples.
Keywords: Particle filters, low & degenerate noise, diffusions, time discretization.
@@
1School of Data Science, The Chinese University of Hong Kong, Shenzhen, Shenzhen, CN.
@@ 1 Introduction
The discrete-time filtering problem considers an unobserved Markov chain $\{X_{n}\}_{n\geq 0}$ (the signal or hidden chain) and associated data $\{Y_{n}\}_{n\geq 1}$. The data at any time $k$ are assumed to depend on $\{X_{n}\}_{n\geq 0}$ only via the position $X_{k}$ of the hidden state. The objective of filtering is to compute expectations with respect to (w.r.t.) the sequence of filtering distributions, i.e. the conditional distribution of the signal at each time $n$, given all the data observed up-to that time, sequentially in the time index. The filtering problem has numerous applications in statistics, applied mathematics, finance and beyond; we refer the reader to the texts [2, 9, 10] and the references therein for a more complete summary.
@@
In most cases of practical interest the filter is unavailable analytically and one must resort to numerical approximations. A plethora of relevant algorithms have been developed, see e.g. [2, 9, 10] for some coverage of the subject. In this article we focus on the approach called particle filters (PFs). These are Monte Carlo, simulation-based methods that simulate a collection of $N\in\mathbb{N}$ samples in parallel and undergo importance sampling and resampling operations. Such methods are very well understood from a mathematical perspective (e.g. [10]) and provide almost surely consistent approximations of the filter as $N\rightarrow\infty$. PFs are typically effective when the dimension of the hidden state, $d_{x}$, is moderate, although there are several recent attempts to further extend the methodology to cover some special-case high-dimensional problems (e.g. [16, 23]).
@@
In this article we focus on the case when the observations have low or degenerate noise. For instance, one could consider the case:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n},$ (1.1)
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}}$, $\Delta\geq 0$, $\{\epsilon_{n}\}_{n\geq 1}$ is a sequence of $\mathbb{R}^{d_{y}}$-valued independent and identically distributed (i.i.d.) noises and $A_{n}$ a sequence of $d_{y}\times d_{x}$ real matrices. of full row rank. For technical reasons that will become apparent in the sequel, we assume that $1\leq d_{y}<d_{x}$ and that $A_{n}$ is of full row rank for every $n\geq 1$. The low noise case could be interpreted as $\Delta$ small (e.g. $\Delta=10^{-6}$) whereas the degenerate case corresponds to $\Delta=0$. As noticed by [30], such a setting is important in many applications where it is expected that the signal dynamics is strongly driving the observed data; see e.g. [3, 14]. From a mathematical perspective, in the case $\Delta=0$ one has to define the filter on the following sequence of manifolds:
$\mathsf{M}^{\star}_{n}=\{x\in\mathbb{R}^{d_{x}}:y_{n}-A_{n}x=0\}.$
@@
Indeed this setting has been explored, for instance, in [30]. In a non-sequential context when one targets a single distribution (e.g., the smoothing law of $X_{0},\dots,X_{n}|y_{1},\dots,y_{n}$ with $n$ fixed), sampling on manifolds has been quite well investigated using Markov chain Monte Carlo (MCMC) in several interesting contexts; see for example [5, 8, 11, 12, 28]. The extension to the filtering problem seems only to be considered in [29, 30], although [29] do not consider a manifold that has been defined by the data such as $\mathsf{M}_{n}$ written above. The approach of [30] is to leverage sequential MCMC and as such it relies on the afore-mentioned MCMC literature. Such methodology is particularly relevant in high-dimensions ($d_{x}$ large) but can be computationally expensive for moderate dimensional problems. Instead, within this latter class of filtering problems it is natural to consider PFs and this is the direction we follow in the present work.
In a more general context, understanding the filtering problem in this low noise regime has been the focus of many studies investigating continuous time dynamics and devising robust (extended) Kalman-Bucy type approximations of the filter [6, 17, 19, 21, 22] or establishing large deviations results for low $\Delta$ [13, 20], see also [18] for a recent review. In general, we need to stress that the filtering distributions are typically analytically intractable except in fairly simple models, such as ones involving linear and Gaussian dynamics. Using approximations such as the extended Kalman filter or variants are biased and most often do not perform well for many models. Given also the increasing accuracy of many sensors used in practice it is important to design PFs that perform well for degenerate or low observation noise case.
The application of plain PFs for low noise problems seldom works well due to the informative likelihoods resulting to very low weights for most particles. This is well-known and illustrated later in this article. The main idea of our work is to reparameterize and cast the given filtering problem as one of sampling on a sequence of manifolds, both in the degenerate setting and in the low noise one. For instance, in the context of (1.1), for $\Delta>0$, we work with the sequence of manifolds:
$\widetilde{\mathsf{M}}_{n}=\{(x,\epsilon)\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:y_{n}-A_{n}x-\Delta^{1/2}\epsilon=0\}.$
A key idea then is that one can carefully define PFs on $\mathsf{M}_{n}$ and $\widetilde{\mathsf{M}}_{n}$, so that the algorithm in the low noise case will be robust by admitting a cost $\mathcal{O}(1)$ w.r.t. $\Delta$, and will thus converge, in a proper sense, to the PF for the degenerate case of $\Delta=0$.
In this paper we explore methodologies with the above properties in the presence of linear observation models such as (1.1). Our contributions can be summarized as follows:
  • (i)
@@
    We derive new PFs for cases of low and degenerate noise, within the setting of linear observation models. We place minimal assumptions on the hidden process $\{X_{n}\}_{n\geq 0}$.
  • (ii)
@@
    In the scenario of (i) and under assumptions, we prove that the $\mathbb{L}_{r}$-error (for a positive integer $r\geq 1$) of the PF in the low noise case will converge, as $\Delta\downarrow 0$, to the $\mathbb{L}_{r}$-error of the PF in the degenerate noise case.
  • (iii)
@@
    For additive Gaussian errors both in the hidden chain and the observations, we derive the optimal proposal and we investigate the corresponding PFs both under low and degenerate noise case. In particular, we show that the weights of the PF for $\Delta>0$ converge to the ones for $\Delta=0$ as $\Delta\downarrow 0$.
  • (iv)
@@
    In the case where the hidden dynamics follow a continuous time diffusion process which is observed at discrete times, we derive a new PF that is robust not only to low noise but also to fine time discretization error. This appears to be the first PF to have this property.
@@
The significance of this work is not only restricted to filtering. The new PFs in this article can be used for parameter estimation associated to diffusion models, such as was done in [1, 15] using the multilevel Monte Carlo method. We note in terms of the fourth contribution, that [7] develop a methodology for the case of hypoelliptic diffusions and degenerate noise (with linear constraints), stating that PFs can be used in their context. However, they do not make the point that this enables the low noise case, which is one of the main findings of this work.
@@ 2.1 State-Space Model
In this section, we work in a general setting in terms of the class of models we consider. Our objective is to present what we call the ‘noise parameterization’ (as an alternative to the ‘natural parametrization’) and the degenerate noise case in full generality. As we stated in the introduction, later on we are only able to develop methodology for linear observation regimes; this is explained in the next section.
We consider the following state-space model (SSM) for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$:
$Y_{n}$ $=H_{n}(X_{n},\epsilon_{n}),$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n}),$
where we have that: $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$; $\epsilon_{n}\in\mathbb{R}^{d_{x}}$, $\nu_{n}\in\mathbb{R}^{d_{y}}$ form sequences of noises that are independent both across $n$ and of each other, with all variables having positive and smooth Lebesgue densities; $H_{n}:\mathbb{R}^{d_{x}+d_{y}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are sequences of continuously differentiable functions with Jacobians that are of full row rank.
@@ 2.2 Natural Parametrization
In the given scenario we can define the filtering densities w.r.t. the $d_{x}$-dimensional Lebesgue measure as follows. Denoting the positive conditional Lebesgue densities of $Y_{n}|X_{n}$ and $X_{n}|X_{n-1}$ as $h_{n}(y_{n}|x_{n})$ and $f_{n}(x_{n}|x_{n-1})$ respectively, we have that the filter (see e.g. [2] and the references therein) at time $1$ writes as follows (with some abuse of notation, here and elsewhere below we use the same symbol both for a distribution and for its density w.r.t. a reference measure implied by the context):
$\pi_{1}(x_{1})\equiv\frac{d\pi_{1}}{d\lambda_{d_{x}}}(x_{1})=\frac{h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathbb{R}^{d_{x}}}h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})\lambda_{d_{x}}(dx_{1})},$ (2.1)
where $\lambda_{d_{x}}$ is the $d_{x}$-dimensional Lebesgue measure and we supress the dependence of $y_{1}$ in the expression for the filtering density $\pi_{1}(x_{1})$. The filter at any subsequent time point $n\geq 1$ has the following prediction-updating formula:
$\pi_{n}(x_{n})\equiv\frac{d\pi_{n}}{d\lambda_{d_{x}}}(x_{n})=\frac{h_{n}(y_{n}|x_{n})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})}{\int_{\mathbb{R}^{d_{x}}}h_{n}(y_{n}|x_{n})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})\lambda_{d_{x}}(dx_{n})}.$ (2.2)
In practice one cannot compute the filter and one must resort to numerical methods. In standard settings, PFs are well known to provide accurate approximations of $\pi_{n}(x_{n})$ and this is for an arbitrary long $n$ under favourable stability assumptions for $F_{n}$ and non-degeneracy of $\epsilon_{n}$. A PF algorithm is in fact designed to target at time $n$ the smoothing density
$\pi_{n}({x}_{1:n})\propto\prod_{k=1}^{n}h_{k}(y_{k}|x_{k})f_{k}({x}_{k}|x_{k-1}),$
from which $\pi_{n}(x_{n})$ can be derived as its final marginal. It should be clear that in this formulation one cannot deal directly with a degenerate case for the conditional likelihood $h_{k}$ and we will redefine an equivalent problem on an appropriate space. In addition, for the low noise case one should note that the region where $h_{k}$ has substantial value will be quite small and becomes a low probability event for traditional sampling proposals to reach it and hence standard PFs can be ineffective. We proceed below with a different parametrization for the same SSM model which will allow for the development of effective numerical methods in the scenario of low noise for the observations.
2.2.1 Particle Filter for the Natural Parameterization
We will present a generic PF for this parameterization when $\Delta>0$. We will use the ancestor representation of the resampling indices denoted as $({a}_{n}^{1},\dots,{a}_{n}^{N})\in\{1,\dots,N\}^{N}$ and we use the convention that ${a}_{0}^{i}=i$ for each $i\in\{1,\dots,N\}$. For the importance sampling part, $q_{n}$ is a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and for the weights we set for any $(x^{\prime},x)\in\mathbb{R}^{2d_{x}}$.
Note that ${q}_{n}$ can depend on $\Delta$ and when $n=1$ we may (or may not) condition on $x_{0}$ in ${q}_{1}$ but write ${q}_{1}(\cdot|x_{0})$ and use $x_{0}^{i}=x_{0}$ for all $i$.
Let $\text{Cat}(\cdot;k,\mathbf{p})$ denote the categorical distribution with $k>0$ categories and $\mathbf{p}=(p_{1},\ldots,p_{k})$ consist a vector of event probabilies such that $p_{i}\geq 0$ and $\sum_{j=1}^{k}p_{j}=1$. Then the PF algorithm proceeds as follows:
  • Initialize: For $i\in\{1,\dots,N\}$, sample $X_{1}^{i}|x_{0}$ from $q_{1}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}\sim\text{Cat}(\cdot;N,\mathbf{w}_{n-1})$ with
    $\mathbf{w}_{n-1}=\left(\frac{w_{n-1}(x_{n-2}^{a_{n-2}^{1}},x_{n-1}^{1})}{\sum_{s=1}^{N}w_{n-1}(x_{n-2}^{a_{n-2}^{s}},x_{n-1}^{s})},\dots,\frac{w_{n-1}(x_{n-2}^{a_{n-2}^{N}},x_{n-1}^{N})}{\sum_{s=1}^{N}w_{n-1}(x_{n-2}^{a_{n-2}^{s}},x_{n-1}^{s})}\right).$
The optimal proposal in terms of minimizing the variance of the weights conditional on ${x}_{n-1}$ is
@@ 2.3 Noise Parametrization and Degenerate Noise
2.3.1 Noise Parametrization
We set $c_{n}(x_{n},\epsilon_{n})=y_{n}-H_{n}(x_{n},\epsilon_{n})$ and define the manifold:
$\widetilde{\mathsf{M}}_{n}=\{(x_{n},\epsilon_{n})\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:c_{n}(x_{n},\epsilon_{n})=0\}.$ (2.3)
Conditionally on $x_{n-1}$ (but not on $y_{n}$), the variables $(x_{n},\epsilon_{n})$ have a Lebesgue density:
$p_{n}(x_{n},\epsilon_{n})=f_{n}(x_{n}|x_{n-1})\,p_{n}(\epsilon_{n}).$
Here $p_{n}(\epsilon_{n})$ is the Lebesgue density of $\epsilon_{n}$. It is useful to think of $p_{n}(x_{n},\epsilon_{n})$ as the prior density for $(x_{n},\epsilon_{n})$ given $x_{n-1}$ before consideration of $y_{n}$. Once $y_{n}$ is received, $(x_{n},\epsilon_{n})$ will lie on the manifold $\widetilde{\mathsf{M}}_{n}$ in (2.3). In particular, the posterior distribution of $(x_{n},\epsilon_{n})$ is their prior with its support now restricted on $\widetilde{\mathsf{M}}_{n}$ (up-to a constant and Jacobian factor). One can obtain an expression for the density of the posterior upon consideration of the appropriate reference measure on the manifold.